OpenStudy (anonymous):

Find the norm of vector r prime of t when r(t) = t^2i + 2sintj + 2costk

OpenStudy (anonymous):

r'(t)=? r(t) is given

OpenStudy (anonymous):

so start with taking derivative

OpenStudy (anonymous):

then square each component , add them

OpenStudy (anonymous):

take square root

OpenStudy (anonymous):

\[\sqrt(4t^2 + 4\cos^2t + 4\sin^2t)\]

OpenStudy (anonymous):

What's the trig identity for cos and sin?

OpenStudy (anonymous):

cos^2 t + sin^2 t =1

OpenStudy (anonymous):

so does the second half of that just become 8?

OpenStudy (anonymous):

4( 1)=4

OpenStudy (anonymous):

Can you explain? I can never follow this one...

OpenStudy (anonymous):

cos(t)^2+sin(t)^2=1 4(cos(t)^2+sin(t)^2)=4

OpenStudy (anonymous):

Ohhhh.... Is that technically combining like terms?

OpenStudy (anonymous):

I guess

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no problem